OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

y = mx + b 5x - 3y = -6 -3y = -5x - 6 y = 5/3 x + 2 m = 5/3

OpenStudy (anonymous):

Thank you!